OpenStudy (anonymous):

You roll two die. What is the probability that that the sum of the numbers with be a multiple of 5. EXPLAIN.

OpenStudy (anonymous):

Min Sum = 2 Max = 12 Multiples of 5 between that = 5,10. Ways to get a sum of 5 = (1, 4) (2,3) ( 4,1) (3,2) {ASSUMING dice are differnt) 4 ways Ways to get 10 = ( 4,6) (6,4) (5,5) = 3 ways Total Outcomes = 6*6 = 36 outcomes. ThereforeReq PRobability = 4+3/36 = 7/36

OpenStudy (anonymous):

Which way did I miss?